OpenStudy (anonymous):

The lengths of two sides of a triangle are shown below: Side 1: 8x2 - 5x - 2 Side 2: 7x - x2 + 3 The perimeter of the triangle is 4x3 - 3x2 + 2x - 6. Part A: What is the total length of the two sides, 1 and 2, of the triangle? Part B: What is the length of the third side of the triangle? Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer.

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

I believe the equation is: 8x^2 - 5x - 2 + 7x - x^2 + 3 + x = 4x^3 - 3x^2 + 2x - 6

OpenStudy (anonymous):

Then you get -4x^3 + 10x^2 + x + 7?

OpenStudy (anonymous):

I don't think so, let me check.

hero (hero):

Add sides 1 and 2, then subtract that sum from the given perimeter to find side 3.

OpenStudy (anonymous):

Oh yeah, duh! I'm so dumb sometimes.

OpenStudy (anonymous):

8x^2 - 5x - 2 + 7x - x^2 + 3 7x^2 + 2x + 1

OpenStudy (anonymous):

So thats the sum from 1 and 2?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

4x^3 - 3x^2 + 2x - 6 -(7x^2 + 2x + 1)

OpenStudy (anonymous):

You have to distribute the -.

OpenStudy (anonymous):

So 4x^3 - 3x^2 + 2x - 6 - 7x^2 - 2x - 1

OpenStudy (anonymous):

Can you solve it from there?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Alright, combine like terms. 4x^3 = 4x^3 -3x^2 - 7x^2 = 2x - 2x = -6 - 1 =

OpenStudy (anonymous):

ok do you still solve from -6 - 1?

OpenStudy (anonymous):

Yeah...

OpenStudy (anonymous):

Why wouldn't you?

OpenStudy (anonymous):

I dont know im so horrible at math its not even funny

OpenStudy (anonymous):

so its -5

OpenStudy (anonymous):

No, -6 - 1 does not = -5. You are subtracting 1 from -6.

OpenStudy (anonymous):

oh -7

OpenStudy (anonymous):

Correct, so what do you have for the others?

OpenStudy (anonymous):

4x^3 = 4x^3 -3x^2 - 7x^2 = 2x - 2x = -6 - 1 = -7

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

I'm back.

OpenStudy (anonymous):

Have you figured it out?

OpenStudy (anonymous):

Yes i did thank you

OpenStudy (anonymous):

Awesome. :D